Please submit a word file with your answers to the following questions:

1. Find and copy a link that could help you translate DNA codons to amino acids.

2. Write the DNA codons and below this the amino acids that correspond to the codons:

AUGUCUCUCACCAACAAGAACGUC

3. For a population with 200 individuals of "AA" genotype, 200 individuals of "Aa" genotype and 600 individuals of "aa" genotype calculate

a. f(A) = p = ?

b. and f(a) = q =?,

c. and examine if the population is in HWE or if there are heterozygote or homozygote genotypes in excess.

4. In a population of bees we sequenced 100 individuals, and found two polymorphic sites. The genotypes found were as follows:

AABB 36

AaBb 48

aabb 16

Please answer

a. Are there missing genotypes?

b. Based on HWE, is there evidence for selection?

c. How can you explain the results (e.g. missing genotypes)?
Answer #1

2. AUGUCUCUCACCAACAAGAACGUC

AUG - Methionine UCU - serine CUC - leucine

ACC - threonine AAC - asparagine AAG - lysine

AAC - asparagine GUC - valine

3. p^2 + 2pq + q^2 = 1 ( according to Hardy Weinberg principle)

a) q^2 = 600/1000 = 0.6 => q = √0.6

f(a)= q = 0.7  

b) Also , p+ q = 1

So, f(A) = p = 1 - 0.7 = 0.3

c) p^2 + 2pq + q^2 = (0.3)^2 + 2×0.3× 0.7 + (0.7)^2 = 1

Therefore , the population is in HWE.